# Fourier123: One Image to High-Quality 3D Object Generation with Hybrid Fourier Score Distillation

### v1

@@ 6 Conclusion and Limitation
In this work, we present Fourier123, a 3D object generation framework that achieves high-quality image-to-3D generation. Two key contributions of our work are: 1) We propose a 2D-3D hybrid Fourier score distillation function, which attempts to optimize 3D representation in both spatial and frequency domains, thereby improving generation quality. 2) Using our hy-FSD, we design a generative Gaussian splatting pipeline, called Fourier123, to achieve efficient 3D generation. It can produce ready-to-use 3D assets from a single image within one minute. We believe that this method of distilling 3D assets from a frequency domain perspective provides a novel approach for the 3D generation task, and this work demonstrates its effectiveness.
Limitation. We share a common problem with existing 3D generation methods: occasional generation failures. We believe this is caused by the inherent randomness of generation task. Fortunately, it is possible to be addressed by repeated generation with different random seeds, since our method only needs 1 minute to generate a 3D asset.
Appendix A Frequency Visualization
@@
In Fig. 6, we display more visual results from the frequency domain to illustrate the frequency differences between the outputs of Zero-1-to-3 (Zero123) and Stable Diffusion (SD). We provide the corresponding frequency results of Zero123 (“Zero123-F”) and SD (“SD-F”) on the right, and their frequency difference (“F-Difference”) at the right end. One can see that in the spatial domain, novel views generated by SD exhibit higher quality compared with those of Zero123, but their structure and identity features are offset. Whilst in the frequency domain, as shown in “F-Difference” of Fig. 6, results of Zero123 and SD exhibit different frequency distributions. We believe that the frequency domain results of SD are more consistent with the frequency distribution of high-quality images, thus using SD to provide frequency supervision. Meanwhile, as Zero123 can produce cross-view geometric-consistent views at the pixel level, we employ Zero123 to supervise our 3D assets in the spatial domain. This combination of distillation is called hy-FSD.
Appendix B Other 2D Appearance Supervision in Frequency Domain
@@
Figure 8: Results of training with super-resolution methods. We alternating the SD model used in hy-FSD with other super-resolution methods, i.e., distilling with Zero123 in spatial domain and using SR methods in frequency domain. One can see that supervising with SR methods in the frequency domain leads to worse results than training with the proposed hy-FSD.
@@
Appendix C Ablation on Fourier123
@@
Settings w/ universal w/ specific
@@
Appendix D Ablation on the Textual Prompt
@@
Appendix E More Visual Results and Comparison

### v2

@@ 6 Conclusion
In this work, we present Fourier123, a 3D object generation framework that achieves high-quality image-to-3D generation. Two key contributions of our work are: 1) We propose a 2D-3D hybrid Fourier score distillation function, which attempts to fully unleash generation ability of T2I model for efficient image-to-3D generation. 2) We design a generative Gaussian splatting pipeline, called Fourier123. It can produce ready-to-use 3D assets from a single image within one minute. We believe that this method of distilling 3D assets from a frequency domain perspective provides a novel approach for the 3D generation task, and this work demonstrates its effectiveness.
Appendix A More Subjective Comparison
The subjective comparison mentioned in Sec. 5.3 is simplified due to the limited space. In Fig. 6, we report the full comparison, including results of the two inference-only methods (LGM [44] and InstantMesh [46]). Furthermore, considering that Fig. 6 only reports the cases that are collected from Internet, that is, lacking the lateral Ground Truth, we also provide the results produced on GSO data from different methods in Fig. 7. Since GSO contains 3D models, we can render lateral Ground Truth for subjective comparison. One can see that the results produced by our method not only contain the best appearances and structures, but are also more consistent with the Ground Truth, which effectively proves the effectiveness of our method.
Figure 8: Frequency difference. We visualize the results generated by Zero123 and SD, and their Amplitude and Phase components in the frequency domain. Zero123 and SD produce very different appearances, with different frequency domain distributions. Overall, the results of SD exhibit better appearances. The amplitude difference between SD and Zero123 is clear that mainly focuses on middle frequency, however, their phase differences are irregular and meaningless.
Appendix B Frequency Visualization
@@
In Fig. 8, we display more visual results from the frequency domain to illustrate the frequency differences between the outputs of Zero-1-to-3 (Zero123) and Stable Diffusion (SD). We provide the corresponding amplitude component of Zero123 (“Zero123-A”) and SD (“SD-A”), the phase component of Zero123 (“Zero123-P”) and SD (“SD-P”), and their difference, i.e., “A-Difference” and “P-Difference” at the right end. One can see that in the spatial domain, novel views generated by SD exhibit higher quality compared with those of Zero123, but their structure and identity features are offset. Whilst in the frequency domain, as shown in “A-Difference” of Fig. 8, results of Zero123 and SD exhibit different amplitude distributions. We believe that the amplitude in frequency domain results of SD is more consistent with that of high-quality images, thus using the frequency amplitude of SD for finer textures. Meanwhile, “P-Difference” shows that the phase components of Zero123 and SD are very different. Considering that phase component represents the content structure of the image, and Zero123 produces cross-view geometric-consistent views at the pixel level but SD cannot, we do not use the frequency phase of SD, but employ Zero123 at the pixel level. This combination of distillation is called hy-FSD.
Appendix C Other 2D Appearance Supervision in Frequency Domain
@@
Figure 10: Results of training with super-resolution methods. We alternate the SD model used in hy-FSD with other super-resolution methods, i.e., distilling with Zero123 in spatial domain and using SR methods in frequency domain. One can see that supervising with SR methods in the frequency domain leads to worse results than training with the proposed hy-FSD.
@@
Table 4: Ablation on Initialization. Fourier123 can still be effective with sphere initialization. The best and the second best results are highlighted in red and blue respectively.
Methods InstantMesh Zero-1-to-3 Magic123 DreamGaussian Fourier123 w/ Sphere Fourier123 w/ LGM
PSNR $\uparrow$ 14.8517 14.8944 14.5827 16.4898 16.7452 21.5049
SSIM $\uparrow$ 0.7942 0.8314 0.7704 0.8231 0.8091 0.8650
LPIPS $\downarrow$ 0.2386 0.1832 0.2764 0.2113 0.2012 0.1112
Appendix D Ablation on Initialization
To prove that the effectiveness of our method is not highly rely on LGM initialization, we use sphere initialization in Fourier123 and compare its performance with others in Tab. 4, results are also calculated on the 100 3D objects from the GSO data. Obviously, this setting still produces competitive results, that is, Fourier123 w/ Sphere outperforms 3 well-known SOTA methods: DreamGaussian, InstantMesh, and Magic123. Considering that DreamGaussian represents methods that using 3D SDS and Magic123 is the SOTA approach that uses both of 2D and 3D SDS. We believe that this comparison can effectively prove the superiority of our method. More importantly, notice that DreamGaussian has to extract mesh from the generated Gaussians and refine mesh texture to get acceptable effects. But Fourier123 can be directly optimized from shpere initialization and get better 3D assets without the second stage fine-tune used in DreamGaussian, which is much better and more efficient.
@@ Appendix D Ablation on Initialization
Appendix E Ablation Score Function on Fourier123
@@
Settings w/ universal text w/ specific texts
@@
Appendix F Ablation on the Textual Prompt
@@
Figure 13: Using supervisions from other feature domains. On the one hand, we apply position embedding to results of SD rather than Fourier transform, building “2D-PE&3D-SDS”. On the other hand, we only use the phase component of results from Zero123 to train. building “2D-FSD&3D-Phase”. Neither of them produce 3D objects well, proving the suitability of our method.
Appendix G Exploring Other Feature Domains
1) Considering that phase component means content structure and we want to use structure priors of Zero123. In the main paper, we directly use its RGB results and build “2D-FSD&3D-SDS” to optimize 3D Gaussian. Here we supplement a study that uses the phase component of Zero123 and discard its RGB results, that it, “2D-FSD&3D-Phase”. 2) On the other hand, we want to demonstrate that the choice of frequency domain is suitable and it cannot be replaced by other feature domains. To this end, for results of SD, we attempt to apply sin-cos Position Embedding to increase its feature channels and take this feature map to optimize. The final optimization function is “2D-PE&3D-SDS”.
Some results of the above mentioned two settings are given in Fig. 13. If altering the Fourier transform with position embedding, results of SD cannot provide texture features effectively. Meanwhile, if we do not use the RGB results of Zero123 but take its phase components, due to the lack of pixel-level supervision, the produced 3D Gaussians cannot convergence well. Consequently, the choice used in main paper is suitable and meaningful.
@@ Appendix G Exploring Other Feature Domains
Appendix H More Visual Results and Comparison
@@
Moreover, the 3D Gaussians produced by our method can be extracted into meshes. Although the quality of extracted meshes is slightly degraded compared to the original 3D Gaussians, they are still exquisite. We provide some meshes in the “meshes” sub-folder of the supplementary materials, which can be visualized by existing 3D softwares such as MeshLab or Blender.
Appendix I Limitation
Due to the inherent randomness of generation methods, we share a common problem with existing 3D generation methods: occasional generation failures. It is possible to be addressed by repeated generation with different random seeds as our method only takes 1 minute to generate once.